Explain what is add-on interest for auto loan, and why a borrower is paying far more...

  1. Explain what is add-on interest for auto loan, and why a borrower is paying far more than the stated rate in this case.
Answer #1
1] When an add-on loan is taken, the interest on the principal
is calculated at the beginning of the term and is then added
to the principal. Thus, the interest forms part of the principal.
The installment amount is then worked out by dividing the
inflated principal by the number of installments.
The formulae that are used to find the installments at the
beginning of the loan period are:
Interest = Principal*Annual interest rate*Term
Installment payments = [Principal+Interest]/Total number of
payments.
Example:
Suppose C takes out an auto loan for $100000, the interest rate
being 4% and the term 4 years, then the workings are done as
below:
Interest = 100000*4%*4 = $         16,000
Monthly installments = (100000+16000)/48 = $           2,417
2] As the interest for the entire term is added to the principal of
the loan at the beginning itself to find the installments, the
actual interest paid by the borrower is much more than the
stated 4%. This is because, interest is not caculated at the
diminishing balance after each installment paid as in a normal
mortgage loan.